OpenStudy (anonymous):

examine the following sets for linear independence:u1= e^x ,u2= e ^−x

OpenStudy (anonymous):

\[u _{1}=e^{x}, u _{2}=e^{-x}, \]

OpenStudy (anonymous):

PLZ HELP

OpenStudy (anonymous):

you know the definition of linear independence?

OpenStudy (anonymous):

u1=e^x, u2=e^-x, are independent if there exist two scalar a,b such that: a(e^x)+b(e^-x)=0

OpenStudy (anonymous):

i know c1u1+c2u2=0; if c1 and c2 is zero then its called linear independence

OpenStudy (anonymous):

multiply both sides by e^x to get: a(e^2x)+b=0 --> a(e^2x)=-b , which only has the solution (a,b)=(0,0).. (a,b)=(0,0) satisfies the definition since 0(e^x)+0(e^-x)=0, and hence they are linearly independent.

OpenStudy (anonymous):

great